OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, 8>, v = <-4, 8> 2.7° -7.3° 5.4° 15.4°

OpenStudy (anonymous):

\[\cos \theta = \frac{ u.v }{ \left| u \right|\left| v \right| }\]

OpenStudy (anonymous):

the cosine of the angle equals the dot product of the two vectors divided by the product of the magnitudes of the two vectors

OpenStudy (phi):

you need to find 3 things: the dot product the "length" of vector u the "length" of vector v can you do that ?

OpenStudy (anonymous):

no i dont knowthe formulas for those

OpenStudy (phi):

you know how to do those, because you did it in the last problem.

OpenStudy (anonymous):

oh ok onesec

OpenStudy (anonymous):

-5*-4+8*8 =84

OpenStudy (phi):

yes that is the dot product. what is the length of each vector ? http://www.dummies.com/how-to/content/calculating-magnitude-with-vectors.html

OpenStudy (anonymous):

sqrt(-5^2+-4^2) = i sqrt(41) sqrt(8^2+8^2) = 8 sqrt(2)

OpenStudy (phi):

try again. First, you use the components of u to find | u | so it is sqrt( (-5)^2 + 8^2 ) also, it is not -(5^2) it is (-5)^2

OpenStudy (anonymous):

sqrt( (-5)^2 + 8^2 ) =89^(1/2) sqrt( (-4)^2 + 8^2 ) = 4 sqrt(5)

OpenStudy (phi):

yes, but I would change them to decimals (keep at least 3 decimals) because the question asks for nearest tenth of a degree. now multiply your lengths you get (about) 9.434 * 8.944= 84.38 now find u dot v / 84.38

OpenStudy (phi):

you figured out u dot v up above. it is 84

OpenStudy (anonymous):

whats u dot v

OpenStudy (anonymous):

84/ 84.38

OpenStudy (phi):

\[ u \cdot v \] is the dot product of u and v you started with the formula \[ \cos x = \frac{ u \cdot v}{|u| \ |v| } \\ \cos x= \frac{84}{84.38}\\ x= \cos^{-1} \left( \frac{84}{84.38}\right)\] can you find x (the angle between the vector u and vector v ) ?

OpenStudy (anonymous):

0.9954965631666271628347949751125859208343209291301256

OpenStudy (phi):

you want the inverse cosine of that number

OpenStudy (phi):

you can type acos(0.995496563) in degrees= into the google search window (or use a calculator)

OpenStudy (anonymous):

5.4 is the answer thx man